62. Methylbenzene (toluene), CHI;CH;(1), is a common

paint thinner. It can be prepared from benzene, Cattal

CH,CI(g) + CoHo(1) > CH;CH;() + HCK(g)

In an investigation, 25.0 g of benzene is combined

with 20.0 g of chloromethane, CH, C1, in the presence

of an appropriate catalyst. (7.4, 7.5)

(a) Calculate the theoretical yield of toluene.

(b) If 22.0 g of CH, CH, is actually obtained, what is the

percentage yield?

(c)

Suggest reasons for the discrepancy between the

actual yield and the theoretical yield.


1
Expert's answer
2022-01-28T07:50:02-0500

"CH_3Cl_{(g)}+C_6H_{6(l)}\to\>C_6H_5CH_{3(l)}+HCl_{(g)}"


Molar mass of CH"_3" Cl "=50.49g\/mol"

C6H6 "=78.11g\/mol"

C6H5CH3 "=92.14g\/mol"



Moles of benzene "=\frac{25}{78.11}=0.3201"


Moles of CH3Cl"=\frac{20}{50.49}=0.3961"


Benzene is the limiting reactant


Moles of Toluene "=" moles of Benzene

"=0.3201"


Theoretical yield "=0.3201\u00d792.14"

"=29.49059"



(b)

Percentage yield "=\frac{22}{29.4905}×100\%"

"=74.6\%"


(c)


Some of the materials;

(1) Were wasted

(2) They failed to react

(3) The product were not captured